OpenStudy (anonymous):

A car accelerates from rest with a 0.2m/^2 acceleration for 5 seconds and kept the same speed for 10 minutes and applied the brakes decelerate at 0.25m/s^2 till it came to a stop. How long is the trip?

OpenStudy (phi):

My first question is are they asking for the length of time or the total distance?

OpenStudy (anonymous):

total distance

OpenStudy (phi):

I would use v = a*t to get the velocity that the car travels at for 10 mins I would use s= 0.5*a*t^2 to get the distance traveled during the accleration then the distance traveled for 10 mins: s= v*t then find the time required to slow to 0 m/s: v= a*t and then use t to find the distance traveled while decelerating.

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (phi):

one catch: velocity will be in m/s so you must change 10 min to 600 sec

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

to find the velocity I would use 0.2m/s^2 for a and 5 seconds for t?

OpenStudy (phi):

yes, \[ v= 0.2 m/s^2 * 5 s= 1.0 m/s \]

OpenStudy (anonymous):

the second one would be 2.5 m for distance right?

OpenStudy (phi):

yes, 2.5 is the distance traveled while accelerating up to 1 m/s now stage 2, find the distance traveled during the next 600 sec

OpenStudy (anonymous):

600m

OpenStudy (phi):

yes

OpenStudy (anonymous):

then what do I add to get the total distance?

OpenStudy (phi):

you need to finish step 3 then find the time required to slow to 0 m/s: v= a*t and then use t to find the distance traveled while decelerating.

OpenStudy (anonymous):

for the time is it 5 or 4 seconds?

OpenStudy (phi):

use decelerate at 0.25m/s^2 and v= 1 m/s in v= a*t to be totally accurate, you are finding the change in velocity \( v_f\) final velocity - \(v_i\) initial velocity= -0.25 m/s^2 * t (it is -0.25 m/s^2 because you are slowing down) of course \(v_f\) is 0 and \(v_i\) is 1 m/s

OpenStudy (anonymous):

so its v= a*t 1= 0.25*t 4 s= t then I add 600m + 4s =604 for the total distance

OpenStudy (phi):

adding meters and seconds ?? remember, in step 1, you traveled 2.5 m while speeding up. use the formula d= 0.5*a*t^2 with a= 0.25 m/s^2 and t= 4 s to find the distance traveled while slowing down then add up the 3 distances.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then is it 602.5?

OpenStudy (phi):

that is the first 2 distances. what about the 3rd distance? (when slowing down)

OpenStudy (anonymous):

its 603.5m

OpenStudy (phi):

no, you must find the 3rd distance using d= 0.5*a*t^2 with a= 0.25 m/s^2 and t= 4 s

OpenStudy (anonymous):

the first distance = 2.5m second= 600m third= 2m 2.5 + 600 +2= 604.5 please say its right

OpenStudy (phi):

yes, it is right

OpenStudy (anonymous):

yes thank soo much you helped me a lot